A chi-square significance test is performed to examine the association between two categorical variables in a 6 x 6 table. What are the degrees of freedom associated with the test statistic? df= 5
Answer #1

As 6*6 table so each of 2 categories has 6 classes. So df will be (6-1)*(6-1) = 5*5 = 25 (Ans)